OpenStudy (anonymous):

Please Help Me?? Use the Law of the Lever to solve the following. Mark sits 4.5 feet from the center of the seesaw to balance a friend who weighs 53 lbs and is sitting 6 feet from the center. How many pounds does Mark weigh? Round to the nearest hundredth. Choose one answer. a. 39.75 lb b. 70.67 lb c. 55.78 lb

OpenStudy (anonymous):

@mrbarry WIll you please please please come help me?

OpenStudy (anonymous):

Technically, this is a physics question, but anyways. Applying the principle that a balanced see-saw has 0 net torue (turning force): Torque = Weight * Shortest distance from fulcrum For friend: 53*6 For Mark: M * 4.5 Since they both are on opposite sides of the see-saw, their torques are equal. => 53 * 6 = 4.5 * M

OpenStudy (anonymous):

So you should get 'b', 70.67.

OpenStudy (anonymous):

Do you see?

OpenStudy (anonymous):

ok thanks. Use the Law of Levers to solve the problem. The number of chairs on a ski lift varies inversely as the distance between them. A ski lift can handle 32 chairs when they are 30 feet apart. If 40 chairs were used, how many feet should be left between each pair? Choose one answer. a. 24 ft b. 37.5 ft c. 32 ft

OpenStudy (anonymous):

@aditkarekatte

OpenStudy (anonymous):

On the see-saw we had: wt * dist = wt * dist here we have: chairs * dist = chairs * dist

OpenStudy (anonymous):

ok..?

OpenStudy (anonymous):

Apply the same principle. Yes, @mrbarry's right

OpenStudy (anonymous):

well duh anything is going to equal itself. i dont get it

OpenStudy (anonymous):

What is mean is that we fill in the facts for one set of chairs on one side and the other set of chairs on the other side. 32 *30 = 40 * dist

OpenStudy (anonymous):

is meant not is mean

OpenStudy (anonymous):

@mrbarry : The number of chairs vary INVERSELY. So it should be 24

OpenStudy (anonymous):

960 = 40x 24 = x

OpenStudy (anonymous):

We are in agreement.

OpenStudy (anonymous):

We are. (Y)

OpenStudy (anonymous):

@kika_rayne ?